Pacific Healthcare is an investor-owned hospital chain. Pacific’s stock is currently selling for $8 per share, current dividend is $0.48 per share (paid on December 31, 2019), and the beta coefficient is 1.2. In addition, you learn that the yield on long-term Treasury bonds is 5.0 percent and that a market return of 11 percent is expected. Most analysts forecast that Pacific will grow about 10 percent per year for the next five years (2020-2024) and 4 percent per year thereafter (2025 and beyond). Based on these growth forecasts, you assembled the forecast dividends per share, as shown in the below exhibit. -

Year/Dividend: 2019/$0.480; 2020/$0.528; 2021/$0.581; 2022/$0.639; 2023/$0.703; 2024/$0.773; 2025/$0.804; 2026/$0.836

Excel Problem (Answers to highlighted cells): What is the estimated value on December 31, 2020, 2021, 2022, 2023, and 2024, assuming the stock is in equilibrium?

Estimated 2020-2024 value of Pacific Healthcare stock:
2021 2022 2023 2024 PV
Year 2-5 dividends $0.58 $0.64 $0.70 $0.77
Year 6+ dividends
Estimated value on December 31, 2020
2022 2023 2024 PV
Year 3-5 dividends $0.64 $0.70 $0.77
Year 6+ dividends
Estimated value on December 31, 2021
Answer #1
Estimated 2020-2024 value of Pacific Healthcare stock:
2020 2021 2022 2023 2024 PV
Year 2-5 dividends $                                  0.58 $                                  0.64 $                                  0.70 $                                      0.77
Present value workings (0.581)/(1+12.2%)^1 = $0.52 (0.639)/(1+12.2%)^2 = $0.51 (0.703)/(1+12.2%)^3 = $0.50 (0.773)/(1+12.2%)^4 = $0.49 $                 2.01
Year 6+ dividends Terminal value = (Dividend x (1+growth rate)) / ( Cost of equity - growth rate) (0.77*1.04)/(12.2%-4%) =$9.77
Present value of Terminal value workings (9.77)/(1+12.2%)^4 = 6.16 6.16
Estimated value on December 31, 2020 2.01+6.16= $8.17
2022 2023 2024 PV
Year 3-5 dividends $                                  0.64 $                                  0.70 $                                      0.77
Present value workings (0.639)/(1+12.2%)^1 = $0.57 (0.703)/(1+12.2%)^2 = $0.56 (0.773)/(1+12.2%)^3 = $0.55 $                 1.68
Year 6+ dividends Terminal value = (Dividend x (1+growth rate)) / ( Cost of equity - growth rate) (0.77*1.04)/(12.2%-4%) =$9.77
Present value of Terminal value workings (9.77)/(1+12.2%)^3 = $6.92 6.92
Estimated value on December 31, 2021 1.68+6.92 = $8.60
Cost of equity = Risk free rate + (Beta x (Market return - Risk free rate))
5%+(1.2*(11%-5%)) = 12.2%
Cost of equity = 12.2%